### v1

@@ 4.2 Line graphs
Let $G$ be a graph and $P=e_{1}e_{2}\ldots e_{k}$ be an isometric path in $L\left(G\right)$. Then $e_{2},e_{3},\ldots,e_{k-1}$ induces an isometric path in $G$.
@@
Let $Q$ be the path induced by the edges $e_{2},e_{3},\ldots,e_{k-1}$ in $G$ and let $u,v$ be the end-vertices of $Q$. If $Q$ is not an isometric path, then another $(u,v)$-path $Q^{\prime}$ in $G$ of length at most $k-3$ exists. Let $e^{\prime}_{1},\ldots,e^{\prime}_{t}$ with $t\leq k-3$ be the edges of $Q^{\prime}$ and notice that $u$ (resp. $v$) is an end-vertex of $e^{\prime}_{1}$ (resp. $e^{\prime}_{k-3}$). Then $e_{2}e^{\prime}_{1}\ldots e^{\prime}_{t}e^{\prime}_{k-1}$ is a shorter path in $L\left(G\right)$, a contradiction. ∎
Lemma 10.
Let $G$ be a graph and $r$ be a vertex such that $ipco\left(\overrightarrow{G_{r}}\right)=k$. For any isometric path $P$, there exists a set $\mathcal{P}$ of $r$-rooted isometric paths with $|\mathcal{P}|\leq k$ such that $|E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)|\leq k$.
Proof.
Since $ipco\left(\overrightarrow{G_{r}}\right)=k$, (by definition) the vertices of $P$ can be covered by $k$ many $r$-rooted isometric paths. Let $\mathcal{P}$ be a set of $r$-rooted isometric paths with $|\mathcal{P}|\leq k$ such that $V(P)\subset\displaystyle\bigcup\limits_{Q\in\mathcal{P}}V(Q)$ and $|E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)|$ is minimized. In other words, $\mathcal{P}$ is a set of $k$ many $r$-rooted isometric paths that cover all vertices of $P$ and cover as many edges of $P$ as possible. We shall show that $\mathcal{P}$ satisfies the lemma statement. Let $X=E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)$ and assume for contradiction that $|X|\geq k+1$. Let the edges of $X$ be ordered as $e_{1},e_{2},\ldots,e_{|X|}$ as they are encountered while traversing $P$ from one end-vertex $a$ to the other end-vertex $b$. For $i\in[|X|]$, let $e_{i}=u_{i}v_{i}$ and while traversing $P$ from $a$ to $b$, we encounter the following ordering of the end-vertices of the edges in $X$: $u_{1},v_{1},u_{2},v_{2},\ldots,u_{|X|},v_{|X|}$. Let $\{i,j\}\subset[|X|]$ and $Q_{i}\in\mathcal{P}$ be a path that contains $u_{i}$. Observe that $Q_{i}$ does not contain $u_{j}$. Otherwise, we could augment $Q_{j}$ to get a set of $r$-rooted isometric paths of cardinality $k$ (different from $\mathcal{P}$), that covers more edges than $\mathcal{P}$. This implies $|\mathcal{P}|\geq k+1$, which is a contradiction. ∎
@@
Let the edges of $P$ be $e_{1},e_{2},\ldots,e_{k}$ and the two end-vertices of $P$ are $u,v$. Then $L\left(P\right)$ has length $k-1$ in $L\left(G\right)$. If $L\left(P\right)$ is not an isometric path, then there exists a $(e_{1},e_{k})$-path in $L\left(G\right)$ that can be written as $e_{1}=e^{\prime}_{1}e^{\prime}_{2}\ldots e^{\prime}_{t}=e_{k}$ such that $t\leq k-2$. Then by Lemma 9, we have that the edges $e^{\prime}_{2},\ldots,e^{\prime}_{t-1}$ induce an isometric path $Q$ in $G$ of length $k-4$. Now, including $e_{1}$ and $e_{k}$ creates a $(u,v)$-path of length at most $k-2$, a contradiction. ∎
@@
Let the edges of $P$ be ordered as $e_{1},e_{2},\ldots,e_{k}$ as they are traversed starting from $u$. We also assume that the end-vertices of $e_{i}$ are $u_{i}$ and $u_{i+1}$ for $i\in[k]$ where $u$ is closer to $u_{i}$ than $u_{i+1}$. Since $L\left(P\right)$ is an isometric path in $L\left(G\right)$ (due to Lemma 11), if $e_{1}=e_{0}$, then the statement is true. Now assume $e_{1}\neq e_{0}$. observe that if $e_{0}$ shares an end-vertex with any edge $e_{j}$ with $j\geq 2$, then $P$ is not an isometric path in $G$. Now, assume that $i$ is the minimum index such that the path induced by $\{e_{0},e_{1},\ldots e_{i}\}$ is not an isometric path in $L\left(G\right)$. From previous arguments, it follows that $i\geq 3$.
Let $Q=e_{0}e^{\prime}_{1}\ldots e^{\prime}_{t}e_{i}$ be an $(e_{0},e_{i})$-isometric path in $L\left(G\right)$. Observe that $t\leq i-2$ and by Lemma 9, the path induced in $G$ by the end-vertices of the edges $e^{\prime}_{1},\ldots,e^{\prime}_{t}$ is an isometric path $P^{\prime}$ in $G$. If the common vertex of $e^{\prime}_{t}$ and $e_{i}$ is $u_{i+1}$, then by appending $u$ with $P^{\prime}$, it is possible to construct an $(u,u_{i+1})$-isometric path of length at most $i-1$ in $G$, which contradicts that $P$ is an isometric path. Hence, the common vertex of $e^{\prime}_{t}$ and $e_{i}$ is $u_{i}$. Let $Q^{\prime}$ denote the path $e_{0}e^{\prime}_{1}\ldots e^{\prime}_{t}e_{i}e_{i+1}\ldots e_{k}$ in $L\left(G\right)$ and observe that the length of $Q^{\prime}$ is at most $k-1$.
We argue that $Q^{\prime}$ is an isometric path in $L\left(G\right)$. If not, then there exists another isometric path $Q^{\prime\prime}=e_{0}e^{\prime\prime}_{1}\ldots e^{\prime\prime}_{x}e_{k}$ in $L\left(G\right)$ such that $x\leq k-3$. Due to Lemma 9, we know that the end-vertices of the edges $e^{\prime\prime}_{1},\ldots,e^{\prime\prime}_{x}$ induce an isometric path $P^{\prime\prime}$ in $G$. By including $u$ and $u_{k+1}$ to $P^{\prime\prime}$ it is possible to construct a $(u,u_{k+1})$-path in $G$ whose length is at most $k-1$. This is a contradiction. ∎
@@
Let $G$ be a graph with $ipco\left(G\right)=k$. Then $ipco\left(L\left(G\right)\right)\leq 3k+2$.
@@
Let $r$ be a vertex of $G$ such that $ipco\left(\overrightarrow{G_{r}}\right)=ipco\left(G\right)=k$. Let $e$ be any edge incident to $r$ and $Q=e_{1}e_{2}\ldots e_{x}$ be an isometric path in $L\left(G\right)$. Due to Lemma 9, $e_{2},e_{3},\ldots e_{x-1}$ induces an isometric path $P$ in $G$. Due to Lemma 10, there exists a set $\mathcal{P}$ of $r$ rooted isometric paths such that $|\mathcal{P}|\leq k$ and $|E(P)\setminus\displaystyle\bigcup\limits_{P^{\prime}\in\mathcal{P}}E(P^{\prime})|\leq k$. Let $e^{\prime}_{1},e^{\prime}_{2},\ldots e^{\prime}_{t}$ with $t\leq k$ are the edges of $P$ that do not belong to any path in $\mathcal{P}$. Due to Lemma 11, for each $P^{\prime}\in\mathcal{P}$, $L\left(P^{\prime}\right)$ is an isometric path in $L\left(G\right)$ and it can be decomposed into two $e$-rooted isometric paths, say $X(P^{\prime})$ and $Y(P^{\prime})$, in $L\left(G\right)$. Now consider the set $\mathcal{Q}=\displaystyle\bigcup\limits_{P^{\prime}\in\mathcal{P}}X(P^{\prime})\cup Y(P^{\prime})$. Notice that, it is possible that none of $E=\{e_{1},e^{\prime}_{1},\ldots,e^{\prime}_{t},e_{x}\}$ may lie in any path of $\mathcal{Q}$. For each $f\in E$, let $Z(f)$ denote an $(e,f)$-isometric path in $L\left(G\right)$. Clearly the set $\mathcal{Q}^{\prime}=\mathcal{Q}\displaystyle\bigcup\limits_{f\in E}Z(f)$ contains only $e$-rooted isometric paths, $|\mathcal{Q}^{\prime}|\leq 3k+2$ and any vertex of $Q$ lies in some isometric path in $\mathcal{Q}^{\prime}$. ∎
Similar arguments as above imply the following.
@@
Let $G$ be a graph with $sipco\left(G\right)=k$. Then $sipco\left(L\left(G\right)\right)\leq 3k+2$.
@@
For $t\geq 1$, the graph $F_{t}$ is defined in three steps as follows: (A) consider a path $P_{t}$ of length $12t-$ 1 and name the vertices $b_{1},c_{1},\ldots,b_{t},c_{t}$ where $b_{1}$ and $c_{t}$ are the two end-vertices of $P_{t}$; (B) consider a new vertex $r$ and for each $1\leq i\leq t$, introduce a new path $Q_{i}$ of length $t-1$ between $r$ and $b_{i}$; (C) introduce a new vertex $r_{0}$ adjacent to $r$. There are no other edges in $F_{t}$ except the ones mentioned above. Sometimes we shall call the paths $Q_{1},Q_{2},\ldots,Q_{t}$ pillars. Now construct $G_{t}$ by considering two disjoint copies (say $C^{1}$ and $C^{2}$) of $F_{t}$ and identify the vertices with label $r_{0}$.
We claim that $ipco\left(G_{t}\right)$ is $t$. Let $a\in V(G)$ be any vertex of $G_{t}$, and without loss of generality, assume that $a\in V(C^{1})$. Consider the copy of the path $P_{t}$ in $C^{2}$ and call it $S$. Observe that $S$ is an isometric path in $G_{t}$ and for any $\{i,j\}\subseteq[t]$, $\mathsf{d}\left(a,b_{i}\right)=\mathsf{d}\left(a,b_{j}\right)$. Moreover, for any $\{i,j\}\subseteq[t]$ there is no $(a,b_{i})$-isometric path in $G_{t}$ that contains $b_{j}$. Therefore $ipco\left(G_{t}\right)\geq t$.
Conversely, for $i\in\{1,2\}$, let $v^{i}$ denote the copy of the vertex $v\in V(F_{t})$ in $C^{i}$. (Recall that $C^{i}$ is isomorphic to $F_{t}$.) Let $T$ be any isometric path in $G_{t}$. If $T$ does not contain the vertex $r_{0}$, it must lie completely in $C^{1}$ or completely in $C^{2}$. Without loss of generality, assume $T$ lies in $C^{1}$. Then the vertices of $T$ can be covered by the set: $\{(r_{0},c^{1}_{j})-\text{isometric paths}\colon j\in[t]\}$. Now consider the case when $T$ contains $r_{0}$. Let $T^{i}=V(T)\cap V(C^{i})$. Observe that $T^{i}$ does not contain vertices from two different pillars. In this case, all the vertices of $T^{i}$ can be covered with an $(r_{0},c^{1}_{j})$-isometric path for some $j\in[t]$. Hence, the vertices of $T$ can be covered with two $r_{0}$-rooted isometric paths in $G_{t}$. Hence $ipco\left(G_{t}\right)\leq t$.
Now we claim that $ipco\left(L\left(G_{t}\right)\right)$ is $12t-$ 1. Observe that $L\left(G_{t}\right)$ is constructed by taking $L\left(C^{1}\right)$ and $L\left(C^{2}\right)$ and identifying the vertices corresponding to the edges $e_{0}=rr_{0}$. For $i\in\{1,2\}$, $L\left(C^{i}\right)$ can be described as follows.
@@
Fix an $i\in\{1,2\}$. Since the copy of $P_{t}$ in $C^{i}$ is an isometric path in $G_{t}$, due to Lemma 11, $S_{i}$ is an isometric path in $L\left(G_{t}\right)$. Consider a pillar $Q_{j}$, $j\in[t]$, in $C^{i}$. Observe that $f_{j}$ (which corresponds to the edge $b_{j}b^{\prime}_{j}$ in $G_{t}$) is at distance $t$ from $e_{0}$ in $L\left(G_{t}\right)$. For $j=1$, $e_{j}$ is at distance $t+1$ from $e_{0}$ and for $j\in[2,t]$, both $e_{j},e^{\prime}_{j}$ are at distance $t+1$ from $e_{0}$. ∎
Using the above claim, we complete the proof. Let $e\in V(L\left(G_{t}\right))$ and without loss of generality assume $e\in V(L\left(C_{1}\right))$. Let $S$ denote the copy of the path $L\left(P_{t}\right)$ in $L\left(C_{2}\right)$. Due to 5.16, $S$ is an isometric path in $L\left(G_{t}\right)$ and all vertices of $L\left(P_{t}\right)$ are equidistant from the vertex $e_{0}=rr_{0}$ in $L\left(G_{t}\right)$. Moreover, there is no isometric path in $L\left(G_{t}\right)$ that contains two vertices of $S$. Therefore, at least $12t-$ 1 many $e$-rooted isometric paths are required to cover the vertices of $S$. Hence $ipco\left(L\left(G_{t}\right)\right)\geq 2t-1$.
Conversely, let $T$ be an isometric path in $L\left(G_{t}\right)$. If $T$ does not contain the vertex $e_{0}$, then the vertices of $T$ lie completely in $L\left(C_{1}\right)$ or completely in $L\left(C_{2}\right)$. In either case, the vertices of $T$ can be covered with $12t-$ 1 many $e_{0}$-rooted isometric paths. If $T$ contains the vertex $e_{0}$, then $T$ can be covered by two $e_{0}$-rooted isometric paths in $L\left(G_{t}\right)$. Hence $ipco\left(L\left(G_{t}\right)\right)\leq 2t-1$. ∎
The proof of Theorem 4(b) follows from Lemmas 13, 14 and 15.

### v2

@@ 4.2 Line graphs
Let $G$ be a graph and $P=e_{1}e_{2}\ldots e_{k}$ with $k\geq 2$ be an isometric path in $L\left(G\right)$. Then $e_{2},e_{3},\ldots,e_{k-1}$ induces an isometric path in $G$.
@@
Let $Q$ be the path induced by the edges $e_{2},e_{3},\ldots,e_{k-1}$ in $G$ and let $u,v$ be the end-vertices of $Q$. If $Q$ is not an isometric path, then another $(u,v)$-path $Q^{\prime}$ in $G$ of length at most k3k-3 exists. Let $e^{\prime}_{1},\ldots,e^{\prime}_{t}$ with $t\leq k-3$ be the edges of $Q^{\prime}$ and notice that $u$ (resp. $v$) is an end-vertex of $e^{\prime}_{1}$ (resp. $e^{\prime}_{t}$). Then $e_{2}e^{\prime}_{1}\ldots e^{\prime}_{t}e_{k-1}$ is a shorter path in $L\left(G\right)$, a contradiction. ∎
@@
Let the edges of $P$ be $e_{1},e_{2},\ldots,e_{k}$ and the two end-vertices of $P$ are $u,v$. Then $L\left(P\right)$ has length k1k-1 in $L\left(G\right)$. If $L\left(P\right)$ is not an isometric path, then there exists a $(e_{1},e_{k})$-path in $L\left(G\right)$ that can be written as $e_{1}=e^{\prime}_{1}e^{\prime}_{2}\ldots e^{\prime}_{t}=e_{k}$ such that $t\leq k-2$. Then by Lemma˜22, we have that the edges $e^{\prime}_{2},\ldots,e^{\prime}_{t-1}$ induce an isometric path $Q$ in $G$ of length k4k-4. Now, including $e_{1}$ and $e_{k}$ creates a $(u,v)$-path of length at most k2k-2, a contradiction. ∎
@@
Let the edges of $P$ be ordered as $e_{1},e_{2},\ldots,e_{k}$ as they are traversed starting from $u$. We also assume that the end-vertices of $e_{i}$ are $u_{i}$ and $u_{i+1}$ for $i\in[k]$ where $u$ is closer to $u_{i}$ than $u_{i+1}$. Since $L\left(P\right)$ is an isometric path in $L\left(G\right)$ (due to Lemma˜23), if $e_{1}=e_{0}$, then the statement is true. Now assume $e_{1}\neq e_{0}$. Observe that if $e_{0}$ shares an end-vertex with any edge $e_{j}$ with $j\geq 2$, then $P$ is not an isometric path in $G$. Now, assume that $i$ is the minimum index such that the path induced by $\{e_{0},e_{1},\ldots e_{i}\}$ is not an isometric path in $L\left(G\right)$. From previous arguments, it follows that $i\geq 3$.
Let $Q=e_{0}e^{\prime}_{1}\ldots e^{\prime}_{t}e_{i}$ be an $(e_{0},e_{i})$-isometric path in $L\left(G\right)$. Observe that $t\leq i-2$ and by Lemma˜22, the path induced in $G$ by the end-vertices of the edges $e^{\prime}_{1},\ldots,e^{\prime}_{t}$ is an isometric path $P^{\prime}$ in $G$. If the common vertex of $e^{\prime}_{t}$ and $e_{i}$ is $u_{i+1}$, then by appending $u$ with $P^{\prime}$, it is possible to construct an $(u,u_{i+1})$-isometric path of length at most i1i-1 in $G$, which contradicts that $P$ is an isometric path. Hence, the common vertex of $e^{\prime}_{t}$ and $e_{i}$ is $u_{i}$. Let $Q^{\prime}$ denote the path $e_{0}e^{\prime}_{1}\ldots e^{\prime}_{t}e_{i}e_{i+1}\ldots e_{k}$ in $L\left(G\right)$ and observe that the length of $Q^{\prime}$ is at most k1k-1.
We argue that $Q^{\prime}$ is an isometric path in $L\left(G\right)$. If not, then there exists another isometric path $Q^{\prime\prime}=e_{0}e^{\prime\prime}_{1}\ldots e^{\prime\prime}_{x}e_{k}$ in $L\left(G\right)$ such that $x\leq k-3$. Due to Lemma˜22, we know that the end-vertices of the edges $e^{\prime\prime}_{1},\ldots,e^{\prime\prime}_{x}$ induce an isometric path $P^{\prime\prime}$ in $G$. By including $u$ and $u_{k+1}$ to $P^{\prime\prime}$ it is possible to construct a $(u,u_{k+1})$-path in $G$ whose length is at most k1k-1. This is a contradiction.
Finally, define $P^{\prime}=e_{0}e_{1}\ldots,e_{i-1}$. The definition of $i$ implies that $P^{\prime}$ is an isometric path in $L\left(G\right)$. Observe that $P^{\prime}$ and $Q^{\prime}$ are two $e_{0}$-rooted isometric paths in $L\left(G\right)$ that cover $L\left(P\right)$. ∎
@@
Let $G$ be a graph and $r$ be a vertex such that $ipco\left(G,r\right)=k$. For any isometric path $P$, there exists a set $\mathcal{P}$ of $r$-rooted isometric paths with $|\mathcal{P}|\leq k$ such that $|E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)|\leq k-1$.
@@
Since $ipco\left(G,r\right)=k$, (by definition) the vertices of $P$ can be covered by at most $k$ many $r$-rooted isometric paths. Let $\mathcal{P}$ be a set of $r$-rooted isometric paths with $|\mathcal{P}|\leq k$ such that $V(P)\subsetneq\displaystyle\bigcup\limits_{Q\in\mathcal{P}}V(Q)$ and $|E(P)\setminus\displaystyle\bigcup\limits_{Q\in\mathcal{P}}E(Q)|$ is minimized. In other words, $\mathcal{P}$ is a set of $k$ many $r$-rooted isometric paths that cover all vertices of $P$ and cover as many edges of $P$ as possible. Observe that, for each $Q\in\mathcal{P}$, the vertices $V(Q)\cap V(P)$ induces a subpath of $P$. Since $|\mathcal{P}|\leq k$, the lemma follows. ∎
@@
Let $G$ be a graph and $r\in V(G)$ be a vertex with $ipco\left(G,r\right)=k$. Then, for any edge $e\in E(G)$ incident on $r$, $ipco\left(L\left(G\right),e\right)\leq 3k+1$.
Proof.
Let $Q=e_{1}e_{2}\ldots e_{x}$ be an isometric path in $L\left(G\right)$. Due to Lemma˜22, $e_{2},e_{3},\ldots e_{x-1}$ induces an isometric path $P$ in $G$. Due to Lemma˜25, there exists a set $\mathcal{P}$ of $r$ rooted isometric paths in $G$ such that $|\mathcal{P}|\leq k$ and $|E(P)\setminus\displaystyle\bigcup\limits_{P^{\prime}\in\mathcal{P}}E(P^{\prime})|\leq k-1$. Let $e^{\prime}_{1},e^{\prime}_{2},\ldots e^{\prime}_{t}$ with $t\leq k-1$ are the edges of $P$ that do not belong to any path in $\mathcal{P}$. Due to Lemmas˜23 and 24, for each $P^{\prime}\in\mathcal{P}$, $L\left(P^{\prime}\right)$ is an isometric path in $L\left(G\right)$ and it can be decomposed into two $e$-rooted isometric paths, say $X(P^{\prime})$ and $Y(P^{\prime})$, in $L\left(G\right)$. Now consider the set $\mathcal{Q}=\displaystyle\bigcup\limits_{P^{\prime}\in\mathcal{P}}X(P^{\prime})\cup Y(P^{\prime})$. Notice that, it is possible that none of $E=\{e_{1},e^{\prime}_{1},\ldots,e^{\prime}_{t},e_{x}\}$ may lie in any path of $\mathcal{Q}$. For each $f\in E$, let $Z(f)$ denote an $(e,f)$-isometric path in $L\left(G\right)$. Clearly the set $\mathcal{Q}^{\prime}=\mathcal{Q}\displaystyle\bigcup\limits_{f\in E}Z(f)$ contains only $e$-rooted isometric paths, $|\mathcal{Q}^{\prime}|\leq 3k+1$ and any vertex of $Q$ lies in some isometric path in $\mathcal{Q}^{\prime}$. ∎
Lemma˜26 implies the following.
Lemma 27.
Let $G$ be a graph. Then $ipco\left(L\left(G\right)\right)\leq 3\cdot ipco\left(G\right)+1$ and $sipco\left(L\left(G\right)\right)\leq 3\cdot sipco\left(G\right)+1$.
@@
For $t\geq 1$, the graph $F_{t}$ is defined in three steps as follows: (A) consider a path $P_{t}$ of length 2t12t-1 and name the vertices $b_{1},c_{1},\ldots,b_{t},c_{t}$ where $b_{1}$ and $c_{t}$ are the two end-vertices of $P_{t}$; (B) consider a new vertex $r$ and for each $1\leq i\leq t$, introduce a new path $Q_{i}$ of length t1t-1 between $r$ and $b_{i}$; (C) introduce a new vertex $r_{0}$ adjacent to $r$. There are no other edges in $F_{t}$ except the ones mentioned above. Sometimes we shall call the paths $Q_{1},Q_{2},\ldots,Q_{t}$ pillars. Now construct $G_{t}$ by considering two disjoint copies (say $C^{1}$ and $C^{2}$) of $F_{t}$ and identify the vertices with label $r_{0}$.
We claim that $ipco\left(G_{t}\right)$ is $t$. Let $a\in V(G)$ be any vertex of $G_{t}$, and without loss of generality, assume that $a\in V(C^{1})$. Consider the copy of the path $P_{t}$ in $C^{2}$ and call it $S$. Observe that $S$ is an isometric path in $G_{t}$ and for any $\{i,j\}\subseteq[t]$, $\mathsf{d}\left(a,b_{i}\right)=\mathsf{d}\left(a,b_{j}\right)$. Hence, for any $\{i,j\}\subseteq[t]$ there is no $(a,b_{i})$-isometric path in $G_{t}$ that contains $b_{j}$. Therefore $ipco\left(G_{t}\right)\geq t$.
Conversely, for $i\in\{1,2\}$, let $v^{i}$ denote the copy of the vertex $v\in V(F_{t})$ in $C^{i}$. (Recall that $C^{i}$ is isomorphic to $F_{t}$.) Let $T$ be any isometric path in $G_{t}$. If $T$ does not contain the vertex $r_{0}$, it must lie completely in $C^{1}$ or completely in $C^{2}$. Without loss of generality, assume $T$ lies in $C^{1}$. For $j\in[t]$, let $S_{j}$ denote a $(r_{0},c^{1}_{j})$-isometric path and let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{t}\}$. Then the vertices of $T$ can be covered by the paths in $\mathcal{S}$. Now consider the case when $T$ contains $r_{0}$ and observe that the vertices of $T$ can be covered by two $r_{0}$-rooted isometric paths in $G_{t}$. Hence $ipco\left(G_{t}\right)\leq t$.
Now we claim that $ipco\left(L\left(G_{t}\right)\right)$ is 2t12t-1. Observe that $L\left(G_{t}\right)$ is constructed by taking $L\left(C^{1}\right)$ and $L\left(C^{2}\right)$ and adding an edge between the vertices corresponding to the edges $e_{0}=rr_{0}$. For $i\in\{1,2\}$, $L\left(C^{i}\right)$ can be described as follows.
@@
Fix an $i\in\{1,2\}$. Since the copy of $P_{t}$ in $C^{i}$ is an isometric path in $G_{t}$, due to Lemma˜23, $S_{i}$ is an isometric path in $L\left(G_{t}\right)$. Consider a pillar $Q_{j}$, $j\in[t]$, in $C^{i}$. Observe that $f_{j}$ (which corresponds to the edge $b_{j}b^{\prime}_{j}$ in $G_{t}$) is at distance $t$ from $e_{0}$ in $L\left(G_{t}\right)$. For $j=1$, $e_{j}$ is at distance $t+1$ from $e_{0}$ and for $j\in[2,t]$, both $e_{j},e^{\prime}_{j}$ are at distance $t+1$ from $e_{0}$. ∎
Using the above claim, we complete the proof. Let $e\in V(L\left(G_{t}\right))$ and without loss of generality assume $e\in V(L\left(C_{1}\right))$. Let $S$ denote the copy of the path $L\left(P_{t}\right)$ in $L\left(C_{2}\right)$. Due to ˜4.2.P, $S$ is an isometric path in $L\left(G_{t}\right)$ and all vertices of $L\left(P_{t}\right)$ are equidistant from the vertex $e_{0}=rr_{0}$ in $L\left(G_{t}\right)$. Moreover, there is no isometric path in $L\left(G_{t}\right)$ that contains two vertices of $S$. Therefore, at least 2t12t-1 many $e$-rooted isometric paths are required to cover the vertices of $S$. Hence $ipco\left(L\left(G_{t}\right)\right)\geq 2t-1$.
Conversely, let $T$ be an isometric path in $L\left(G_{t}\right)$. If $T$ does not contain the vertex $e_{0}$, then the vertices of $T$ lie completely in $L\left(C_{1}\right)$ or completely in $L\left(C_{2}\right)$. In either case, the vertices of $T$ can be covered by 2t12t-1 many $e_{0}$-rooted isometric paths. If $T$ contains the vertex $e_{0}$, then $T$ can be covered by two $e_{0}$-rooted isometric paths in $L\left(G_{t}\right)$. Hence $ipco\left(L\left(G_{t}\right)\right)\leq 2t-1$. ∎
The proof of Theorem˜8(b) follows from Lemmas˜26, 27 and 28.